OpenStudy (anonymous):

How do I solve this? Write an equation of a line that passes through the point (0, 4) and whose slope is 4. Answer y = x + 4 y = 4x + 4 y = x – 4 y = x

OpenStudy (anonymous):

If its a line you know it will be of the form $$y=mx+b$$ Where the slope is m

OpenStudy (anonymous):

Sense you know its slope and you have a point on the line, you can solve for b. Thus you will have m and b, and you can completely determine an equation for representing the line.

OpenStudy (anonymous):

$$y=4x+b$$ $$4=4\times 0 +b$$ $$4=b$$

OpenStudy (anonymous):

So we have $$y=4x+4$$ And its the second one.